# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 5.2. Validation
To improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or worker-based checks (Figure 6).
@@
Title. Validation, to improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or worker-based checks. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses), Constraints (increase expense and latency), Accuracy (check for correctness), Creativity (confirm elements of quality exist), Task specific (can validate specific qualities). Section 2: tactics used. Workers (LLMs, crowdworkers, user), Subtasks (evaluate, transform), Architectures (sequential, redundant, communicative).
@@
Motivation. Individual workers can give low quality responses. Unaddressed, these can lead to cascading errors through the workflow (Wu et al., 2022a; Goto et al., 2016; Little et al., 2010b; Bernstein et al., 2010; Little et al., 2010a). Therefore, higher quality subtask outcomes throughout the workflow can have myriad benefits on the overall outcome quality for both accuracy and creativity focused tasks, increasing robustness to low-quality worker responses (Ambati et al., 2012; Bernstein et al., 2010).
@@
Implementation. The validation process is often sequential, in which content is first generated before being validated (Gero et al., 2023). In parallel and iterative redundant workflows, errors from prior passes can be found and fixed or rejected (Zaidan and Callison-Burch, 2011; Hahn et al., 2016) and differing opinions can be merged (Kittur et al., 2011; Luther et al., 2015) through transform and evaluate subtasks (Wu et al., 2023a, c; Zharikova et al., 2023; Kulkarni et al., 2012; Ambati et al., 2012; Schick et al., 2022). Users can break ties (Pietrowicz et al., 2013) and intervene if there are cascading errors (Kulkarni et al., 2012). Non-worker methods, such as taking the majority vote of the outputs of redundant evaluate subtasks can also be effective. Automated methods of flagging malicious workers, gold-standard validation steps, or requiring a certain agreement threshold among workers within a redundant architecture can also provide validation (Lasecki et al., 2015; Parameswaran et al., 2023; Huffaker et al., 2020; Liem et al., 2011).
@@
Debate-based communicative workflows can promote reflection on disagreements (Kobayashi et al., 2018; Chen et al., 2019; Du et al., 2023; Liang et al., 2023). Using different models and roles in debate shows promise (Du et al., 2023), as an LLM judge in a debate was found to more likely agree with its same underlying model (Liang et al., 2023). Other workflows use adaptable architectures to make all production always editable so workers can fix errors in a adaptable manner (Liu et al., 2018; LaToza et al., 2014).
@@
Consequences. Costs increase when validation involves more calls to workers. The literature shows mixed results for an LLM’s ability to validate its own responses in evaluate subtasks. Some studies suggest that adding a self-validation step improves outcomes (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023), while others suggest that LLMs are likely to agree with their previous outputs, even when incorrect, and propose debate-based communicative workflows as an alternative (Liang et al., 2023; Du et al., 2023).
@@
Examples. Examples of the validation strategy include: the “Verify” step of Soylent (Bernstein et al., 2010), a Dual-Pathways structure for transcription (Liem et al., 2011), and self-evaluation guided reasoning (Xie et al., 2023).
@@ 5.3. User guidance
To help the user explore the design space and align creative outcomes to their vision, incorporate their input into the workflow at runtime (Figure 7).
@@
Title. User Guidance. To help the user explore the design space and align creative outcomes to their vision, incorporate their input into the workflow at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Users learn through interaction), Constraints (Increased effort), Accuracy (Review uncertainties), Creativity (Guide the creative vision), Task general (Can predefine less before runtime). Section 2: tactics used. Workers (user), Subtasks (evaluate, focus), Architectures (redundant, dynamic, communicative).
@@
Motivation. Users may want to use the workflow to reduce the effort of exploring of the design space, to align the workflow’s output to their creative vision, or some combination thereof (Nebeling et al., 2016; Willett et al., 2012; Kulkarni et al., 2014; Park et al., 2022). If a process includes data that is inherently uncertain it could benefit from more explicit guidance by the user (Chang et al., 2017; Willett et al., 2012; Cheng and Bernstein, 2015).
@@
Implementation. This strategy incorporates the user into a critical role within the workflow. Users can guide outcomes by dynamically answering questions and providing feedback (Salehi et al., 2017) or monitor dynamic workflows to catch cascading errors (Kulkarni et al., 2012) with evaluate or transform subtasks. Users can be called upon when necessary with a “panic button” that other workers can invoke (Kulkarni et al., 2012), or automatically when ties occur (Pietrowicz et al., 2013). With a shared communicative interface between users and all workers, users can intervene when needed and make real-time suggestions for prototyping directions (Nebeling et al., 2016; Lasecki et al., 2015). Users may review at intermediate steps to edit or regenerate sections of the workflow in evaluate or focus subtasks (Park et al., 2022; Kim et al., 2023b; Zhang et al., 2023a; Kulkarni et al., 2014). A redundant iterative architecture may also include multiple iterations with user revisions between each one (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Kim et al., 2023b). Sometimes disagreements raised with redundant architectures may not need to be resolved within the workflow and can instead support an explorable interface for users to directly manipulate (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019). They can sort and filter the returned data (Willett et al., 2012; Cheng and Bernstein, 2015; Mohanty et al., 2019), use a slider to control output characteristics like text length (Bernstein et al., 2010), and consider category labels at multiple levels of precision (Chang et al., 2017).
@@
Crowdsourcing workflows integrate the user in more varied ways than LLM chaining, including providing results without aggregation that the user can sort and filter (Lin et al., 2014a; Mohanty et al., 2019), user guidance for prototyping (Cheng and Bernstein, 2015; Kulkarni et al., 2014; Retelny et al., 2014), and worker-initiated user revision (Salehi et al., 2017).
@@
Secondary benefits. Incorporating the user into the process at runtime can reduce the amount of pre-definition of workflow steps needed, making workflows more generally applicable (Nebeling et al., 2016; Kulkarni et al., 2014).
@@
Examples. Examples of the user guidance strategy include: Second Opinion (Mohanty et al., 2019), Social Simulacra (Park et al., 2022), and WearWrite (Nebeling et al., 2016).
@@ 5.4. Adaptable architectures
Title. Adaptable Architectures. To build workflows that require less effort to design and that generalize to many inputs, use adaptive architectures that build themselves at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Cascading errors and infinite loops, Less transparency and consistency), Constraints (Reduce expense and latency), Creativity (Adapt to edge cases), Task general (Adaptable for task generality, Request tasks with no pre-definition, Decomposable input required). Section 2: tactics used. Subtasks (focus, partition), Architectures (branching, dynamic, communicative).
@@
Motivation. It is challenging and requires high effort to design a rigid and pre-defined workflow that can effectively generalize to many tasks or that can accommodate for many edge-cases, especially in creative tasks (Kim and Monroy-Hernandez, 2016; Valentine et al., 2017; Kittur et al., 2011; Kulkarni et al., 2012). Instead, designers may want workflows that are adaptable and responsive (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Wu et al., 2023a; Nair et al., 2023b; Creswell and Shanahan, 2022) and to support a variety of inputs without the effort of explicit pre-definition (Kittur et al., 2011; Kulkarni et al., 2012; Creswell et al., 2022; Creswell and Shanahan, 2022).
@@
Implementation. Adaptable workflows are not rigid in their progression of subtasks but instead are flexible and generate parts of the workflow architecture at runtime (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Kittur et al., 2011; Huang et al., 2023c, b). Workflows can incorporate subtasks defined by the crowd with partition and focus subtasks in dynamic and communicative architectures (Ahmad et al., 2011; Zhang et al., 2012; Agapie et al., 2015; Retelny et al., 2014; Reppert et al., 2023). Changes in these adaptable workflows may be made by the worker (LaToza et al., 2014; Valentine et al., 2017; Retelny et al., 2014), user (Retelny et al., 2014; Nebeling et al., 2016; Valentine et al., 2017; Kulkarni et al., 2012), or an algorithm (Bragg et al., 2013; Chen et al., 2019; Dai et al., 2013; Goto et al., 2016). Although branching architectures are predefined, they also allow for adaptability in which different inputs take different paths.
@@
Consequences. However, there are tradeoffs. Adaptable workflows can lead to poor quality worker-created subtasks and instructions, causing cascading errors and infinite loops (Kulkarni et al., 2012). They also are restricted to tasks with an easily decomposable input (Kittur et al., 2011).
@@
Chaining papers highly emphasize the importance of transparency (Creswell et al., 2022; Creswell and Shanahan, 2022; Reppert et al., 2023; Wu et al., 2022a; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2021, 2023a), increasing the risk of using adaptable workflows which may cause unforeseeable errors as each step is not predefined. The degree to which adaptable workflows degrade this transparency is yet to be explicitly explored.
@@
Some effects of adaptable workflows on worker experiences mattered for crowdsourcing, but do not matter for chaining. Crowdsourcing work has allowed workers to chose their own tasks as an incentive to improve task performance (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014), but the need to incentivize workers does not apply to chaining. One crowdsourcing study that found users were frustrated with too much adaptability, an emotion that does not apply to LLMs (Valentine et al., 2017).
@@
Title. Worker recruitment, To improve outcome quality at each step, curate a high quality worker pool by filtering workers by their skills and improving worker capabilities over time. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increase cost and effort of filtering). Section 2: tactics used. Workers (LLMs, crowdworkers).
@@
Examples. Examples of adaptable architectures include: map-reduce (Kittur et al., 2011), Flash Teams (Retelny et al., 2014), and microtask programming (LaToza et al., 2014).
@@
5.5. Worker recruitment
To improve outcome quality at each step, curate a high quality worker pool by filtering workers by their skills and improving worker capabilities over time (Figure 9).
Motivation. As workers power the workflow, workers that can provide higher-quality responses can help support outcome quality (Bragg et al., 2018; Liu et al., 2016).
Implementation. In crowdsourcing, recruiting high-quality workers can mean filtering crowdworkers by pre-specified qualifications or using platforms with more specialized workers such as Upwork (Valentine et al., 2017). For LLM chains, model choice and finetuning impact worker abilities (Cobbe et al., 2021; Bursztyn et al., 2022; Creswell et al., 2022; Creswell and Shanahan, 2022).
Consequences. Higher quality workers may be more expensive (Christoforaki and Ipeirotis, 2014; Huang et al., 2015; Bursztyn et al., 2022). There is also effort and expense involved in filtering crowdworkers for skill (Kulkarni et al., 2014; Zaidan and Callison-Burch, 2011; Valentine et al., 2017) or finetuning models to improve specific skills (Bursztyn et al., 2022).
Examples. Examples that pay particular attention to worker recruitment include Flash Organizations (Valentine et al., 2017), and chains that finetune models (Bursztyn et al., 2022) or use multiple models (Liang et al., 2023).
Related strategies. Worker recruitment works hand in hand with aligning subtasks to worker capabilities. Worker recruitment must also be considered when using adaptive architectures to define what kind of workers to hire during runtime.

### v4

@@ 5.2. Validation
To improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or actor-based checks (Figure 7).
@@
Title. Validation, to improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or actor-based checks. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses), Constraints (increase expense and latency), Objectivity (check for correctness), Subjectivity (confirm elements of quality exist), Task specific (can validate specific qualities). Section 2: tactics used. Actors (LLMs, crowdworkers, user), Subtasks (evaluate, improve, merge), Architectures (sequential, redundant, communicative).
@@
Motivation. Individual actors can give low quality responses. Unaddressed, these can lead to cascading errors through the workflow [C(Little et al., 2010a, b; Goto et al., 2016; Bernstein et al., 2010) L(Wu et al., 2022a)]. Therefore, higher quality subtask outcomes throughout the workflow can have myriad benefits on the overall outcome quality for both objectivity and subjectivity focused tasks, increasing robustness to low-quality actor outputs [C(Ambati et al., 2012; Bernstein et al., 2010)].
@@
Implementation. The validation process is often sequential, in which content is first generated before being validated [L(Gero et al., 2023)]. In parallel and iterative redundant workflows, errors from prior passes can be found and fixed or rejected through improve subtasks [C(Hahn et al., 2016; Zaidan and Callison-Burch, 2011)] and differing opinions can be merged [C(Luther et al., 2015; Kittur et al., 2011)] through merge and evaluate subtasks [C(Ambati et al., 2012; Kulkarni et al., 2012) L(Wu et al., 2023a; Zharikova et al., 2023; Wu et al., 2023b; Schick et al., 2022)]. Users can break ties [C(Pietrowicz et al., 2013)] and intervene if there are cascading errors [C(Kulkarni et al., 2012)]. Non-worker methods, such as taking the majority vote of the outputs of redundant evaluate subtasks can also be effective. Automated methods of flagging inattentive crowdworkers, gold-standard validation steps, or requiring a certain agreement threshold among actors within a redundant architecture can also provide validation [C(Liem et al., 2011; Huffaker et al., 2020; Lasecki et al., 2015) L(Parameswaran et al., 2023)].
@@
Debate-based communicative workflows can promote reflection on disagreements [C(Kobayashi et al., 2018; Chen et al., 2019) L(Liang et al., 2023; Du et al., 2023)]. Using different models and roles in debate shows promise [L(Du et al., 2023)], as an LLM judge in a debate was found to more likely agree with its same underlying model [L(Liang et al., 2023)]. Other workflows use adaptable architectures to make all production always editable so actors can fix errors in an adaptable manner [C(Liu et al., 2018; LaToza et al., 2014)].
@@
Consequences. Costs increase when validation involves more calls to actors. The literature shows mixed results for an LLM’s ability to validate its own responses in evaluate subtasks. Some studies suggest that adding a self-validation step improves outcomes [L(Madaan et al., 2024; Gero et al., 2023; Yao et al., 2024; Xie et al., 2023)], while others suggest that LLMs are likely to agree with their previous outputs, even when incorrect, and propose debate communicative workflows as an alternative [L(Du et al., 2023; Liang et al., 2023)].
@@
Examples. Examples of the validation strategy include: the “Verify” step of Soylent [C(Bernstein et al., 2010)], a Dual-Pathways structure for transcription [C(Liem et al., 2011)], and self-evaluation guided reasoning [L(Xie et al., 2023)].
@@ 5.3. User guidance
To help the user explore the design space and align subjective outcomes to their vision, incorporate their input into the workflow at runtime (Figure 8).
@@
Title. User Guidance. To help the user explore the design space and align subjective outcomes to their vision, incorporate their input into the workflow at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Users learn through interaction), Constraints (Increased effort), Objectivity (Review uncertainties), Subjectivity (Guide the creative vision), Task general (Can predefine less before runtime). Section 2: tactics used. Actors (user), Subtasks (evaluate, focus), Architectures (redundant, dynamic, communicative).
@@
Motivation. Users may want to use the workflow to reduce the effort of exploring the design space, to align the workflow’s output to their creative vision, or some combination thereof [C(Kulkarni et al., 2014; Nebeling et al., 2016; Willett et al., 2012) L(Park et al., 2022)]. If a process includes inherently uncertain data it could benefit from more explicit guidance by the user [C(Willett et al., 2012; Cheng and Bernstein, 2015; Chang et al., 2017)].
@@
Implementation. This strategy incorporates the user into a critical role within the workflow. Users can guide outcomes by dynamically answering questions and providing feedback [C(Salehi et al., 2017)] or monitor dynamic workflows to catch cascading errors [C(Kulkarni et al., 2012)] with evaluate subtasks. Users can be called upon when necessary with a “panic button” that other workers can invoke [C(Kulkarni et al., 2012)], or automatically when ties occur [C(Pietrowicz et al., 2013)]. With a shared communicative interface between users and all workers, users can intervene when needed and make real-time suggestions for prototyping directions [C(Nebeling et al., 2016; Lasecki et al., 2015)]. Users may review at intermediate steps to edit or regenerate sections of the workflow in evaluate or focus subtasks [C(Kulkarni et al., 2014) L(Park et al., 2022; Kim et al., 2023b; Zhang et al., 2023a)]. A redundant iterative architecture may also include multiple iterations with user revisions between each one [C(Kulkarni et al., 2014; Cheng and Bernstein, 2015; Willett et al., 2012) L(Kim et al., 2023b; Park et al., 2022)]. Sometimes disagreements raised with redundant architectures may not need to be resolved within the workflow and can instead support an explorable interface for users to directly manipulate [C(Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019; Willett et al., 2012)]. They can sort and filter the returned data [C(Cheng and Bernstein, 2015; Mohanty et al., 2019; Willett et al., 2012)], use a slider to control output characteristics like text length [C(Bernstein et al., 2010)], and consider category labels at multiple levels of precision [C(Chang et al., 2017)].
@@
Crowdsourcing workflows integrate the user in more varied ways than LLM chaining, including providing results without aggregation to sort and filter [C(Mohanty et al., 2019; Lin et al., 2014a)], user guidance for prototyping [C(Kulkarni et al., 2014; Cheng and Bernstein, 2015; Retelny et al., 2014)], and actor-initiated user revision [C(Salehi et al., 2017)].
@@
Secondary benefits. Incorporating the user into the process at runtime can reduce the amount of pre-definition of workflow steps needed, making workflows more generally applicable [C(Kulkarni et al., 2014; Nebeling et al., 2016)].
@@
Examples. Examples of the user guidance strategy include: Second Opinion [C(Mohanty et al., 2019)], Social Simulacra [L(Park et al., 2022)], and WearWrite [C(Nebeling et al., 2016)].
@@ 5.4. Adaptable architectures
Title. Adaptable Architectures. To build workflows that require less effort to design and that generalize to many inputs, use adaptive architectures that build themselves at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Cascading errors and infinite loops, Less transparency and consistency), Constraints (Reduce expense and latency), Subjectivity (Adapt to edge cases), Task general (Adaptable for task generality, Request tasks with no pre-definition, Decomposable input required). Section 2: tactics used. Subtasks (focus, partition), Architectures (branching, dynamic, communicative).
@@
Motivation. It is challenging and requires high effort to design a rigid and pre-defined workflow that can effectively generalize to many tasks or that can accommodate for many edge-cases, especially in subjective tasks [C(Kulkarni et al., 2012; Kittur et al., 2011; Valentine et al., 2017; Kim and Monroy-Hernandez, 2016)]. Instead, designers may want workflows that are adaptable and responsive [C(Kulkarni et al., 2014; Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Chen et al., 2019) L(Wu et al., 2023a; Creswell and Shanahan, 2022; Nair et al., 2023b)] and to support a variety of inputs without the effort of explicit pre-definition [C(Kulkarni et al., 2012; Kittur et al., 2011) L(Creswell et al., 2022; Creswell and Shanahan, 2022)].
@@
Implementation. Adaptable workflows are not rigid in their progression of subtasks but instead are flexible and generate parts of the workflow architecture at runtime [C(Kulkarni et al., 2014; Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kittur et al., 2011; Chen et al., 2019) L(Huang et al., 2023b, c)]. Workflows can incorporate subtasks defined by the crowd with partition and focus subtasks in dynamic and communicative architectures [C(Agapie et al., 2015; Retelny et al., 2014; Zhang et al., 2012; Ahmad et al., 2011) L(Reppert et al., 2023)]. Changes in these adaptable workflows may be made by the worker [C(Valentine et al., 2017; LaToza et al., 2014; Retelny et al., 2014)], user [C(Kulkarni et al., 2012; Nebeling et al., 2016; Valentine et al., 2017; Retelny et al., 2014)], or an algorithm [C(Bragg et al., 2013; Dai et al., 2013; Chen et al., 2019; Goto et al., 2016)]. Although branching architectures are predefined, they also allow for adaptability in which different inputs take different paths.
@@
Consequences. However, there are tradeoffs. Adaptable workflows can lead to poor quality worker-created subtasks and instructions, causing cascading errors and infinite loops [C(Kulkarni et al., 2012)]. They also are restricted to tasks with an easily decomposable input [C(Kittur et al., 2011)].
@@
Chaining papers highly emphasize the importance of transparency [L(Wu et al., 2023a; Creswell et al., 2022; Creswell and Shanahan, 2022; Wu et al., 2021; Gero et al., 2023; Wu et al., 2022a; Reppert et al., 2023)], increasing the risk of using adaptable workflows which may cause unforeseeable errors as each step is not predefined. The degree to which adaptable workflows degrade this transparency is yet to be explicitly explored.
@@
Some effects of adaptable workflows on worker experiences matter for crowdsourcing, but do not matter for chaining. Crowdsourcing work has allowed crowdworkers to chose their own tasks as an incentive to improve task performance [C(Chilton et al., 2019; Nebeling et al., 2016; LaToza et al., 2014; Mahyar et al., 2018; Kittur et al., 2011; Zhang et al., 2012)], but the need to incentivize does not apply to chaining. One crowdsourcing study found users were frustrated with too much adaptability, an emotion that does not apply to LLMs [C(Valentine et al., 2017)].
@@
Title. Actor selection, To improve outcome quality at each step, curate a high quality actor pool by filtering actors by their skills and improving actor capabilities over time. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increase cost and effort of filtering). Section 2: tactics used. Actors (LLMs, crowdworkers).
@@
Examples. Examples of adaptable architectures include: map-reduce [C(Kittur et al., 2011)], Flash Teams [C(Retelny et al., 2014)], and microtask programming [C(LaToza et al., 2014)].
@@
5.5. Actor selection
To improve outcome quality at each step, curate a high quality actor pool by filtering actors by their skills and improving actor capabilities over time (Figure 10).
Motivation. Actors that provide higher-quality responses support outcome quality [C(Drapeau et al., 2016; Huang et al., 2015; Ahmad et al., 2011)].
Implementation. In crowdsourcing, selecting high-quality actors can mean filtering crowdworkers by pre-specified qualifications or using platforms with more specialized crowdworkers such as Upwork [C(Valentine et al., 2017)]. With support, crowdworkers can improve performance over time [C(Chen et al., 2019; Weir et al., 2015; Drapeau et al., 2016; Kobayashi et al., 2018)]. For LLM chains, model choice, specialization, and finetuning impact abilities [L(Creswell et al., 2022; Cobbe et al., 2021; Bursztyn et al., 2022; Creswell and Shanahan, 2022)].
Consequences. Higher quality actors may be more expensive [C(Huang et al., 2015; Christoforaki and Ipeirotis, 2014) L(Bursztyn et al., 2022)], although smaller specialized models may reduce expense [L(Bursztyn et al., 2022)]. There is also effort and expense involved in filtering crowdworkers for skill [C(Kulkarni et al., 2014; Zaidan and Callison-Burch, 2011; Valentine et al., 2017)] or finetuning models to improve specific skills [L(Bursztyn et al., 2022)].
Examples. Examples that pay particular attention to actor selection include Flash Organizations [C(Valentine et al., 2017)], and chains that finetune models [L(Bursztyn et al., 2022)] or use multiple models [L(Liang et al., 2023)].
Related strategies. Actor selection works hand in hand with aligning subtasks to actor capabilities. Actor selection must also be considered when using adaptive architectures to define what kind of workers to hire during runtime.